OpenStudy (anonymous):

is V25 = ±5 always? why?

OpenStudy (anonymous):

V25 = \[\sqrt{25}\]

OpenStudy (anonymous):

actually it denotes positive square root of 25

OpenStudy (anonymous):

so answer will be 5

OpenStudy (lgbasallote):

@akash123 you're talking about principal square roots...

OpenStudy (anonymous):

Because 5² = (=5)² = 25

OpenStudy (anonymous):

yes it's... but if anyone use this symbol of sqrt then it implies positive sqrt only eg as we write sqrt(sin^2x) = I sinx I

OpenStudy (lgbasallote):

(=5)^2? interesting

OpenStudy (anonymous):

I think it's a convention only and moreover it's frequently used in calculus

OpenStudy (anonymous):

So YES, the even root always yield 2 results :)

OpenStudy (anonymous):

It's basic Pre Algebra!

OpenStudy (anonymous):

Because 5² = (-5)² = 25

OpenStudy (anonymous):

yup. but why? hm..

OpenStudy (anonymous):

x^2= x* x

OpenStudy (anonymous):

25= 5*5 and 25 =(-5)* (-5)

OpenStudy (anonymous):

i know that. i mean... argh. i cant speak english very well. but i think that's enough thanks for your hardwork xD